A particular shoe franchise knows that its stores will not show a profit unless they gross over $940,000 per year. Let A be the event that a new store grosses over $940,000 its first year. Let B be the event that a store grosses over $940,00 its second year. The franchise has an administrative policy of closing a new store if it does not show a profit in either of the first 2 years. The accounting office at the franchise provided the following information: 63% of all the franchise stores show a profit the first year; 67% of all the franchise stores show a profit the second year (this includes stores that did not show a profit the first year); however, 81% of the franchise stores that showed a profit the first year also showed a profit the second year. Compute the following. (Enter your answers to four decimal places.)

(a)    P(A)


(b)    P(B)


(c)    P(B | A)


(d)    P(A and B)


(e)    P(A or B)


(f)    What is the probability that a new store will not be closed after 2 years?


What is the probability that a new store will be closed after 2 years?
Answer #1

A PO-PCture shuss profit for ist year) 6 P(1) - P (store shop forft for and year) = 0.67 9 PCB(A) = P (shows profit for and yiP (new store will be closed) = P (no profit in both year) = 1- 0.5103 = 0.4897